OpenStudy (anonymous):

Use the quadratic formula to find the zeros of the function. Round to tenths if necessary. Write your solution in set brackets with no spaces, such as {-1,2}. y=3x^2+5x

OpenStudy (anonymous):

3x^2+5x = 0 x(3x+5) = 0 x = 0 x = -5/3 So, zeroes of the function are: {0, -5/3}